OpenStudy (anonymous):

A waiter earned $28 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 16 bills.

OpenStudy (lgbasallote):

hint: let x = number of $1 bills y = number of $5 bills

OpenStudy (anonymous):

i ve already tried solving it! i just can figure it out!

OpenStudy (lgbasallote):

the total number of bills is 16...so number of $1 bill + number of $5 bill = 6

OpenStudy (lgbasallote):

that's 16 not 6

OpenStudy (lgbasallote):

now just transform it into algebra

OpenStudy (anonymous):

well you could use trial and error for this one since its not a very large

OpenStudy (anonymous):

soo 2?

OpenStudy (lgbasallote):

wait...what are you supposed to do? find the number of $5 bills? or just create an equation?

OpenStudy (anonymous):

max amount of $5 bills in $28 and then work your way down

OpenStudy (anonymous):

find the number of $5 bills in 28 only sing 16 1 or 5 dollar bills

OpenStudy (anonymous):

using*

OpenStudy (lgbasallote):

okay here's a hint...use this equation (i suggest you understand what it means) x + y = 16 x + 5y = 28 since y is the number of 5 dollar bills...solve for y

OpenStudy (anonymous):

makes no since

OpenStudy (lgbasallote):

x + y = 16 x is the number of 1 dollar bill y is the number of 5 dollar bill so the number of 1 dollar bill + number of 5 dollar bill = 16

OpenStudy (lgbasallote):

the total amount is 28 dollars so it would be $1 times the amount of 1 dollar bill wich is x so 1 times x = x 5 dollars times the amount of 5 dollar bills = 5 times x = 5y so x + 5y = 28

OpenStudy (lgbasallote):

anyway use systems of equations to solve for y x + y = 16 x + 5y = 28

OpenStudy (anonymous):

soo 3 $5 bills

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks